Recent Log Entries

IPM1G03 is missing from LERF ISOTOPE beam spikes screens

Log Number: 3862528Logbook(s): Body: 

Updating HCO for LERF bpms this morning. I did a HCO of the LERF bpms last Thursday but forgot to update the LERF HCO application even tho I did make elog entries Thursday. While doing this I noticed that IPM1G03 was missing from the LERF ISOTOPE beam line spike screens. This happens all the time. This bpm is sometimes used by Ops when first setting up Beam for the Isotope Line. I will send a email to software to get the screens updating again.

COMMENT ON VME crate upgrade to IOCIN3

Log Number: 3862527Logbook(s): Body: 

After opening the back VME panel I found the ribbon connections connected to the backplane of slot #7. This card is labeled Injector HV Controls. There are no connections to the front of the board. The plan is to reconnect the ribbon cables from the old HV switch interface chassis to the backplane of the new crate. Once we check out complete functionality from this IOC, I will be able to determine if the connections in question are still being used by EPICS. Crate swap can continue when ready.

LERF Hot-Standby Operator SIGN IN

Log Number: 3862521Logbook(s): Body:  LERF Hot-Standby Operator SIGN IN FORM LERF Hot-Standby Operator: Frierson
DATE: 01/12/2021
SHIFT: Owl

Operator Comments: Going into Hot Stand by until Day shift. There is an issue with the DLPC that cannot be addressed until then

The above LERF Hot-Standby Operator has received pass down from: Forman

LERF RF Finally Back Up

Log Number: 3862504Logbook(s): Entry Makers: kernsBody: 

1) After replacing a failed IOCFEL1 chassis due to power supply failure that threw all sorts of errors and 2) replacing the cooling fan on Quarter 4 gun that didn't allow it to go to HV and 3) No output at all on Quarter 3 after getting it back into RF required a RFCM replacement and finicky tuning and 4) FEL02 cavity 8 that lost all drive again since Friday's RFCM replacement that lasted an hour before crash. I replaced the RFCM again tonight and lost the drive entirely after only minutes of RF. Turned out to be a dying 5W amplifier that took it's time dying !

CREW CHIEF SIGN IN

Log Number: 3862445Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 01-11-21
Time : 16:33:47
Day : Monday
Shift : Swing

Crew Chief : Moser
Program Deputy: Mounts
Team : C
Arm : RADCON
Designated SSO: Moser
LERF Operator: Forman
Operators : N/A

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

buncher heater ON now

Log Number: 3862429Logbook(s): Entry Makers: poelkerBody: 

Buncher heater power restored from the break (overlooked that one). Setpoint 164 F to make buncher resonant 748.5 MHz

UITF LLRF testing

Log Number: 3862423Logbook(s): Entry Makers: plawskiBody: 

started testing cavity 2 on the booster in sel and gdr using the existing algorithm to make baseline measurement
in sel mode
gradient 8.3 MV/m
discriminator 3 Hz rms, -4 Hz min, 5 Hz max. There is a dominant frequency component at 26.4 Hz. 59 Hz dominant mode from the previous measurements doesn't show up anymore
in gdr mode with FIR
gradient 8.3 MV/m
detune frequency 8 Hz rms, -21 Hz min, 17 Hz max and the dominant mode is still the same.
integrated phase noise from 1 Hz to 3 Mhz is 240 fs
amplitude regulation is .04 % rms

LERF Operator SIGN IN

Log Number: 3862416Logbook(s): Body:  LERF OPERATOR SIGN IN FORM OPERATOR:    Freeman
DATE:             01/11/2021
SHIFT:               Swing

This stamp certifies that the above operator has read/reviewed/received the following:

iocitfrfstp restored

Log Number: 3862407Logbook(s): Tags: Body: restorelib 2.2.1 Target restore date was 2021-01-11 15:45:39 The output from the myRestore execution was as follows: iocitfrfstp Rf12Step@2021-01-11 15:45 120(120) restored system@2021-01-11 15:45 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ued -A -s -i iocitfrfstp (success) Procedure return status: Success

COMMENT ON Trim Utility Chassis Upgrade - All CEBAF Trim Racks

Log Number: 3862401Logbook(s): Body: 

The three newly assembled Utility Chassis field tested in
UITF trim racks namely ITF02, ITF03 and ITF07. Have
been tested remotely ITF02 and ITF03 passed all stages of
the remote testing. ITF07 had some problems with the
individual reset lines. This will not stop UITF from running.
All three UITF trim racks are up and running, with all
magnets set to Zero.

We also tested another upgraded Utility Chassis in both
Local / Remote mode and it pass all stages of the test.
the Utility Chassis will be place out in E1 service building
tomorrow.

Diggs / Kissinger

COMMENT ON VME crate upgrade to IOCIN3

Log Number: 3862371Logbook(s): Body: 

Tony, lets take a quick pause on replacing the VME crate until
we can talk (I will be in at 6am). I found several connections
going to the VME backplane from the old HV Switch Interface
chassis. I believe the connections are no longer active, but I
am not 100% sure. I would like to remove the back panel from
the existing crate and document where each connector goes.
We can reconnect in the same locations to the new crate until I
can figure out if they are still active.

Pages